# Nonasymptotic Analysis of Stochastic Gradient Descent with the Richardson–Romberg Extrapolation

### v2

@@ 4 Richardson-Romberg extrapolation
where we have defined, for $\theta\in\mathbb{R}^{d}$, the following vector-valued functions:
$\begin{split}\psi(\theta)&=(1/2)\nabla^{3}f(\theta^{*})(\theta-\theta^{\star})^{\otimes 2}\,,\\ G(\theta)&=(1/2)\left(\int_{0}^{1}t^{2}\nabla^{4}f(t\theta^{\star}+(1-t)\theta)\,\mathrm{d}t\right)(\theta-\theta^{\star})^{\otimes 3}\,.\end{split}$ (37)
@@
where $B_{3}\in\mathbb{R}^{d}$ is a vector such that $\|B_{3}\|\leq\frac{L}{\mu}\mathsf{C}_{1}+\frac{L\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{5/2}}$, and
@@ Discussion.
Proof of Section 4 is provided in Section D.1. It is important to acknowledge that there are numerous Rosenthal-type inequalities for dependent sequences in the literature. Section 4 can be viewed as an analogue to the classical Rosenthal inequality for strongly mixing sequences, see (Rio, 2017, Theorem 6.3). However, it should be emphasized that the Markov chain $\{\theta_{k}^{(\gamma)}\}_{k\in\mathbb{N}}$ is geometrically ergodic under the assumptions A 1-A 3($p$) only in sense of the weighted Wasserstein semi-metric $\mathbf{W}_{{c}}(\xi,\xi^{\prime})$ with respect to a cost function $c$ defined in (16). As a result, the sequence $\{\theta_{k}^{(\gamma)}\}_{k\in\mathbb{N}}$ does not necessarily satisfy strong mixing conditions. At the same time, $\{\theta_{k}^{(\gamma)}\}_{k\in\mathbb{N}}$ satisfies the $\tau$-mixing condition, see Merlevède et al. (2011). However, the considered function $\psi(\theta)$ is quadratic, which means that the result of (Merlevède et al., 2011, Theorem 1) can not be directly applied. Bounds similar to (41) have been explored in (Durmus et al., 2023), but in Section 4 we obtain the bound with tighter dependence of the right-hand side upon $\gamma$. Below we provide the $p$-th moment bound together with corollary for the step size $\gamma$ optimized w.r.t. $n$.
@@
$\begin{split}\mathsf{E}_{\nu}^{1/p}[\|\operatorname{H}^{\star}(\bar{\theta}_{n}^{(RR)}-\theta^{\star})\|^{p}]&\leq\frac{c_{1}\sqrt{\operatorname{Tr}{\Sigma_{\varepsilon}^{\star}}}p^{1/2}}{n^{1/2}}+\frac{\mathsf{C}_{\operatorname{RR},5}}{n\gamma^{1/2}}+\frac{\mathsf{C}_{\operatorname{RR},6}\gamma^{1/2}}{n^{1/2}}+\mathsf{C}_{\operatorname{RR},7}\gamma^{3/2}\\ &\qquad\qquad+\frac{c_{2}p\tau_{p}}{n^{1-1/p}}+\frac{\mathsf{C}_{\operatorname{RR},8}}{n}+\mathcal{R}_{5}(n,\gamma,\|\theta_{0}-\theta^{\star}\|)\,,\end{split}$ (42)
where $c_{1}=60\mathrm{e}$ and $c_{2}=60$ are absolute constants from the Pinelis version of Rosenthal inequality (Pinelis, 1994, Theorem 4.1), problem-specific constants $\mathsf{C}_{\operatorname{RR},5}$ to $\mathsf{C}_{\operatorname{RR},8}$ are defined in Appendix D (equation (101)), and
@@
Proof of Theorem 3 is provided in Appendix D. Note that the result above is a direct generalization of Theorem 2, which reveals the same scaling of the step size $\gamma$ with respect to $n$. To the best of our knowledge, this is the first analysis of a first-order method, which provides a $p$-th moment bound with $p>2$ and the second-order term of order $\mathcal{O}\bigl{(}n^{-3/4}\bigr{)}$ while keeping the precise leading term related to the minimax-optimal covariance matrix $\Sigma_{\varepsilon}^{\star}$. Such results were previously obtained for the setting of linear stochastic approximation (LSA), see Mou et al. (2020); Durmus et al. (2024). Thus, Richardson-Romberg extrapolation applied to strongly convex minimization problems allows for the restoration of the $p$-th moment error bounds as obtained in the LSA setting.
@@ 5 Numerical results
In this section we illustrate numerically the scale of the second-order terms in equation (40) in Section 4. We show that, for a particular minimization problem, setting $\gamma=n^{-1/2}$, we achieve the scaling of the second-order terms in root-MSE bounds of order $\mathcal{O}(n^{-3/4})$. We consider the problem
@@
$n\|\bar{\theta}_{n}^{(RR)}-\theta^{\star}\|^{2}\text{ and }n^{3/2}\|\bar{\theta}_{n}^{(RR)}+n^{-1}\sum\nolimits_{k=n+1}^{2n}{\varepsilon_{k+1}}(\theta^{\star})-\theta^{\star}\|^{2}\,,$
@@ 6 Conclusion
7 Acknowledgements
The work of M. Sheshukova, A. Naumov, and S. Samsonov was prepared within the framework of the HSE University Basic Research Program. The work of E. Moulines has been partly funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ 7 Acknowledgements
Throughout this appendix we use $c_{0}$ for an absolute constant, which values may vary from line to line. In addition, when the upper index of $\theta_{k}^{(\gamma)}$ or $\theta_{k}^{(2\gamma)}$ is omitted, we assume the result applies to iterations of $\theta_{k}^{(\gamma)}$. The corresponding results for $\theta_{k}^{2\gamma}$ can be obtained by substituting $2\gamma$ instead of $\gamma$. We also provide some additional definitions related to the Markov kernels and kernel couplings, particularly useful when considering convergence in Wasserstein semimetric. Detailed exposition can be found in (Douc et al., 2018, Chapter 20).
Let ${\rm Q}(z,A)$ be a Markov kernel on $(\mathsf{Z},\mathcal{Z})$. We say that ${\rm K}$ is a Markov coupling of ${\rm Q}$ if for all $(z,z^{\prime})\in\mathsf{Z}^{2}$ and $A\in\mathcal{Z}$, ${\rm K}((z,z^{\prime}),A\times\mathsf{Z})={\rm Q}(z,A)$ and ${\rm K}((z,z^{\prime}),\mathsf{Z}\times A)={\rm Q}(z^{\prime},A)$. If ${\rm K}$ is a kernel coupling of ${\rm Q}$, then for all $n\in\mathbb{N}$, ${\rm K}^{n}$ is a kernel coupling of ${\rm Q}^{n}$ and for any $\mathcal{C}\in\mathscr{C}(\xi,\xi^{\prime})$, $\mathcal{C}{\rm K}^{n}$ is a coupling of $(\xi{\rm Q}^{n},\xi^{\prime}{\rm Q}^{n})$. Moreover, it holds that
$\mathbf{W}_{{c}}(\xi{\rm Q}^{n},\xi^{\prime}{\rm Q}^{n})\leq\int_{\mathsf{Z}\times\mathsf{Z}}{\rm K}^{n}c(z,z^{\prime})\mathcal{C}(\mathrm{d}z\mathrm{d}z^{\prime})\,,$ (44)
see (Douc et al., 2018, Corollary 20.1.4). For any probability measure $\mathcal{C}$ on $(\mathsf{Z}^{2},\mathcal{Z}^{\otimes 2})$, we denote by $\mathbb{P}_{\zeta}^{\rm K}$ and $\mathsf{E}_{\zeta}^{\rm K}$ the probability and the expectation on the canonical space $((\mathsf{Z}^{2})^{\mathbb{N}},(\mathcal{Z}^{\otimes 2})^{\otimes\mathbb{N}})$ such that the canonical process $\{(Z_{n},Z_{n}^{\prime}),n\in\mathbb{N}\}$ is a Markov chain with initial probability $\mathcal{C}$ and Markov kernel ${\rm K}$. We write $\mathsf{E}_{z,z^{\prime}}^{{\rm K}}$ instead $\mathsf{E}_{\delta_{z,z^{\prime}}}^{{\rm K}}$.
@@
$\qquad=\mathsf{E}\bigl{[}\|\theta_{k}-\theta^{\star}\|^{2}-2\gamma\langle\nabla F(\theta_{k},\xi_{k+1}),\theta_{k}-\theta^{\star}\rangle+\gamma^{2}\|F(\theta_{k},\xi_{k+1})\|^{2}|\mathcal{F}_{k}\bigr{]}\,.$
@@
$\mathsf{E}[\|\theta_{k+1}-\theta^{\star}\|^{2}|\mathcal{F}_{k}]\leq\|\theta_{k}-\theta^{\star}\|^{2}-2\gamma\langle\nabla f(\theta_{k})-\nabla f(\theta^{\star}),\theta_{k}-\theta^{\star}\rangle+\gamma^{2}\mathsf{E}[\|F(\theta_{k},\xi_{k+1})\|^{2}|\mathcal{F}_{k}]\,.$
@@
$\mathsf{E}[\|F(\theta_{k},\xi_{k+1})\|^{2}|\mathcal{F}_{k}]$ $=\mathsf{E}[\|F(\theta_{k},\xi_{k+1})-F(\theta^{\star},\xi_{k+1})+{\varepsilon_{k+1}}(\theta^{\star})\|^{2}|\mathcal{F}_{k}]$
@@
Note that $2(1-\gamma\mu)^{r/2}\leq 2$ for any $r\leq m(\gamma)-1$ and $22(1-\gamma\mu)^{m(\gamma)/2}\leq 1/$ 2. Hence, applying the result of (Douc et al., 2018, Theorem 20.3.4), we obtain that the Markov kernel ${\rm Q}_{\gamma}$ admits a unique invariant distribution $\pi_{\gamma}$. Applying (44), we get
@@
We first prove (22) and (23) and introduce some additional notations. Under A 1A 3($2$), we define a matrix-valued function $\mathcal{C}(\theta):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ as
@@
where $\bar{\theta}_{\gamma}$ is defined in (22), $\mathcal{C}(\theta)$ is defined in (49), and $B_{1}\in\mathbb{R}^{d}$ satisfies $\|B_{1}\|\leq\frac{\operatorname{L}}{\mu}\mathsf{C}_{1}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{5/2}}$, with
$\mathsf{C}_{1}=\frac{\sqrt{\operatorname{L}}\tau_{2}^{2}}{\sqrt{\mathsf{C}_{\operatorname{step},6}}\mu}+\frac{1}{2}\left(\left(\frac{\operatorname{L}^{2}\mathsf{D}_{\operatorname{last},2}}{\mu^{3/2}}+\frac{\operatorname{L}\sqrt{\mathsf{D}_{\operatorname{last},2}}}{\sqrt{\mu}}\right)\tau_{2}^{2}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{3/2}}+\frac{\operatorname{L}^{1/2}\mathsf{D}_{\operatorname{last},4}^{2}\tau_{4}^{4}}{4\mu^{2}\mathsf{C}_{\operatorname{step},6}^{3/2}}\right)$ (51)
@@
for any matrix $A\in\mathbb{R}^{d\times d}$, and $B_{2}\in\mathbb{R}^{d\times d}$ is a matrix, such that $\|B_{2}\|\leq\mathsf{C}_{1}$.
@@
[…]
@@
where $B\in\mathbb{R}^{d\times d}$ satisfies $\|B\|\leq\mathsf{C}_{2}$. Using (57) together with C 1($6$), we obtain that $\bar{\Sigma}_{\gamma}$ is a solution to the matrix equation
@@
$\|B^{\prime}\|_{{F}}\leq\mathsf{C}_{2}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{3/2}}+\frac{\operatorname{L}^{1/2}\mathsf{D}_{\operatorname{last},4}^{2}\tau_{4}^{4}}{4\mu^{2}\mathsf{C}_{\operatorname{step},6}^{3/2}}\,.$ (59)
@@
$\leq\frac{\sqrt{\gamma}\operatorname{L}\tau_{2}^{2}}{\mu}+\frac{1}{2}\left(\mathsf{C}_{2}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{3/2}}+\frac{\operatorname{L}^{1/2}\mathsf{D}_{\operatorname{last},4}^{2}\tau_{4}^{4}}{4\mu^{2}\mathsf{C}_{\operatorname{step},6}^{3/2}}\right)$
$\leq\frac{\sqrt{\operatorname{L}}\tau_{2}^{2}}{\sqrt{\mathsf{C}_{\operatorname{step},6}}\mu}+\frac{1}{2}\left(\mathsf{C}_{2}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{3/2}}+\frac{L^{1/2}\mathsf{D}_{\operatorname{last},4}^{2}\tau_{4}^{4}}{4\mu^{2}\mathsf{C}_{\operatorname{step},6}^{3/2}}\right)\,,$
@@
$\|\int_{\mathbb{R}^{d}}\mathcal{C}(\theta)\pi_{\gamma}(\mathrm{d}\theta)-\mathcal{C}(\theta^{\star})\|_{{F}}\leq\mathsf{C}_{2}\gamma^{1/2}\,,$
where the constant $\mathsf{C}_{2}$ is given by
$\mathsf{C}_{2}=\left(\frac{\operatorname{L}^{2}\mathsf{D}_{\operatorname{last},2}}{\mu^{3/2}}+\frac{\operatorname{L}\sqrt{\mathsf{D}_{\operatorname{last},2}}}{\sqrt{\mu}}\right)\tau_{2}^{2}\,.$ (62)

### v3

@@ 4 Richardson-Romberg extrapolation
where we have defined the following vector-valued functions:
$\psi(\theta)=\frac{1}{2}\nabla^{3}f(\theta^{*})(\theta-\theta^{\star})^{\otimes 2}\,,\quad G(\theta)=\frac{1}{2}\left(\int_{0}^{1}t^{2}\nabla^{4}f(t\theta^{\star}+(1-t)\theta)\,\mathrm{d}t\right)(\theta-\theta^{\star})^{\otimes 3}\,.$ (38)
@@
where $B_{3}\in\mathbb{R}^{d}$ is a vector such that $\|B_{3}\|\leq\mathsf{C}_{b,1}$, where $\mathsf{C}_{b,1}$ and is constant independent of $\gamma$ and is defined in (LABEL:eq:_def_ConstPRb1,2) and
@@ Discussion.
Proof of Section 4 is provided in Section D.1. It is important to acknowledge that there are numerous Rosenthal-type inequalities for dependent sequences in the literature. Section 4 can be viewed as an analogue of the classical Rosenthal inequality for strongly mixing sequences, see (Rio, 2017, Theorem 6.3). However, it should be emphasized that the Markov chain $\{\theta_{k}^{(\gamma)}\}_{k\in\mathbb{N}}$ is geometrically ergodic under the assumptions A 1-A 3($p$) only in sense of the weighted Wasserstein semi-metric $\mathbf{W}_{{c}}(\xi,\xi^{\prime})$ with cost function $c$ defined in (16). As a result, the sequence $\{\theta_{k}^{(\gamma)}\}_{k\in\mathbb{N}}$ does not necessarily satisfy strong mixing conditions. Bounds similar to (42) have been explored in (Durmus et al., 2023), but in Section 4 we obtain tighter dependence of the right-hand side upon $\gamma$. Below we provide the $p$-th moment bound together with corollary for the step size $\gamma$ optimized w.r.t. $n$.
@@
[…]
where $c_{1}=60\mathrm{e}$ and $c_{2}=60$ are absolute constants from the Pinelis version of Rosenthal inequality (Pinelis, 1994, Theorem 4.1), and problem-specific constants $\mathsf{C}_{\operatorname{RR},5}$ to $\mathsf{C}_{\operatorname{RR},8}$ are defined in Appendix D (equation (100)), and
@@
Proof of Theorem 9 is provided in Appendix D. Note that the result above is a direct generalization of Theorem 6, which reveals the same scaling of the step size $\gamma$ with respect to $n$. To the best of our knowledge, this is the first analysis of a first-order method, which provides a $p$-th moment bound with $p>2$ and the second-order term of order $\mathcal{O}\bigl{(}n^{-3/4}\bigr{)}$ while keeping the precise leading term related to the asymptotically optimal covariance matrix $\Sigma_{\varepsilon}^{\star}$. Such results were previously obtained for the setting of linear stochastic approximation (LSA), see Mou et al. (2020); Durmus et al. (2025). Thus, Richardson-Romberg extrapolation applied to the strongly convex minimization problems allows to mimic the $p$-th moment error bounds that were previously obtained in the LSA setting.
@@ 5 Numerical results
In this section we numerically illustrate the scale of the second-order terms in equation (41) in Section 4. We show that, for a particular minimization problem, setting $\gamma=n^{-1/2}$, we achieve the scaling of the second-order terms in root-MSE bounds of order $\mathcal{O}(n^{-3/4})$. We consider the problem
@@
$n\|\bar{\theta}_{n}^{(RR)}-\theta^{\star}\|^{2}\text{ and }n^{3/2}\|\bar{\theta}_{n}^{(RR)}-\theta^{\star}+n^{-1}\sum\nolimits_{k=n+1}^{2n}{\varepsilon_{k+1}}(\theta^{\star})\|^{2}\,$,
@@ 6 Conclusion
Acknowledgement
The work of M. Sheshukova, D. Belomestny, A. Naumov, and S. Samsonov was prepared within the framework of the HSE University Basic Research Program. The work of E. Moulines has been partly funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Acknowledgement
Throughout this appendix we use $c_{0}$ for an absolute constant, which values may vary from line to line. In addition, when the upper index of $\theta_{k}^{(\gamma)}$ or $\theta_{k}^{(2\gamma)}$ is omitted, we assume the result applies to iterations of $\theta_{k}^{(\gamma)}$. The corresponding results for $\theta_{k}^{(2\gamma)}$ can be obtained by substituting $2\gamma$ instead of $\gamma$. We provide some additional definitions related to the Markov kernels and kernel couplings, particularly useful when considering convergence in Wasserstein semimetric. Detailed exposition can be found in (Douc et al., 2018, Chapter 20).
Let ${\rm Q}(z,A)$ be a Markov kernel on $(\mathsf{Z},\mathcal{Z})$. A Markov kernel $\rm K$} on $(\mathsf{Z}^{2},\mathcal{Z}^{\otimes 2})$ is called a kernel coupling of $({\rm Q},{\rm Q})$ (that is, of $\rm Q$} with itself), if for all $(z,z^{\prime})\in\mathsf{Z}^{2}$ and $A\in\mathcal{Z}$, ${\rm K}((z,z^{\prime}),A\times\mathsf{Z})={\rm Q}(z,A)$ and ${\rm K}((z,z^{\prime}),\mathsf{Z}\times A)={\rm Q}(z^{\prime},A)$. If $\rm K$} is a kernel coupling of $({\rm Q},{\rm Q})$, then for all $n\in\mathbb{N}$, ${\rm K}^{n}$ is a kernel coupling of $({\rm Q}^{n},{\rm Q}^{n})$ and for any $\Pi\in\mathscr{C}(\xi,\xi^{\prime})$, $\Pi{\rm K}^{n}$ is a coupling of $(\xi{\rm Q}^{n},\xi^{\prime}{\rm Q}^{n})$. Moreover, it holds that
$\mathbf{W}_{{c}}(\xi{\rm Q}^{n},\xi^{\prime}{\rm Q}^{n})\leq\int_{\mathsf{Z}\times\mathsf{Z}}{\rm K}^{n}c(z,z^{\prime})\Pi(\mathrm{d}z\mathrm{d}z^{\prime})\,$, (45)
see (Douc et al., 2018, Corollary 20.1.4). For any probability measure $\Pi$ on $(\mathsf{Z}^{2},\mathcal{Z}^{\otimes 2})$, we denote by $\mathbb{P}_{\Pi}^{\rm K}$ and $\mathsf{E}_{\Pi}^{\rm K}$ the probability and the expectation on the canonical space $((\mathsf{Z}^{2})^{\mathbb{N}},(\mathcal{Z}^{\otimes 2})^{\otimes\mathbb{N}})$ such that the canonical process $\{(Z_{n},Z_{n}^{\prime}),n\in\mathbb{N}\}$ is a Markov chain with initial probability $\Pi$ and Markov kernel $\rm K$}. We write $\mathsf{E}_{z,z^{\prime}}^{{\rm K}}$ instead $\mathsf{E}_{\delta_{z,z^{\prime}}}^{{\rm K}}$.
@@
$\qquad=\mathsf{E}\bigl{[}\|\theta_{k}-\theta^{\star}\|^{2}-2\gamma\langle\nabla F(\theta_{k},\xi_{k+1}),\theta_{k}-\theta^{\star}\rangle+\gamma^{2}\|\nabla F(\theta_{k},\xi_{k+1})\|^{2}|\mathcal{F}_{k}\bigr{]}\,.$
@@
$\mathsf{E}[\|\theta_{k+1}-\theta^{\star}\|^{2}|\mathcal{F}_{k}]\leq\|\theta_{k}-\theta^{\star}\|^{2}-2\gamma\langle\nabla f(\theta_{k})-\nabla f(\theta^{\star}),\theta_{k}-\theta^{\star}\rangle+\gamma^{2}\mathsf{E}[\|\nabla F(\theta_{k},\xi_{k+1})\|^{2}|\mathcal{F}_{k}]\,.$
@@
$\mathsf{E}[\|\nabla F(\theta_{k},\xi_{k+1})\|^{2}|\mathcal{F}_{k}]$ $=\mathsf{E}[\|\nabla F(\theta_{k},\xi_{k+1})-\nabla F(\theta^{\star},\xi_{k+1})+{\varepsilon_{k+1}}(\theta^{\star})\|^{2}|\mathcal{F}_{k}]$
@@
Note that $2(1-\gamma\mu)^{r/2}\leq 2$ for any $r\leq m(\gamma)-1$ and $2(1-\gamma\mu)^{m(\gamma)/2}\leq 1/2$. Hence, applying the result of (Douc et al., 2018, Theorem 20.3.4), we obtain that the Markov kernel ${\rm Q}_{\gamma}$ admits a unique invariant distribution $\pi_{\gamma}$. Applying (45), we get
@@
We first prove (22) and (23) and introduce some additional notations. Under A 1A 3(22), we define a matrix-valued function $\mathcal{C}(\theta):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ as
@@
where $\bar{\theta}_{\gamma}$ is defined in (22), $\mathcal{C}(\theta)$ is defined in (50), and $B_{1}\in\mathbb{R}^{d}$ satisfies $\|B_{1}\|\leq\mathsf{C}_{b,1}$, with
$\begin{split}&\mathsf{C}_{b,1}=\frac{\operatorname{L}}{\mu}\mathsf{C}_{b,2}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{5/2}}\\ &\mathsf{C}_{b,2}=\frac{\sqrt{\operatorname{L}}\tau_{2}^{2}}{\sqrt{\mathsf{C}_{\operatorname{step},6}}\mu}+\frac{1}{2}\left(\left(\frac{\operatorname{L}^{2}\mathsf{D}_{\operatorname{last},2}}{\mu^{3/2}}+\frac{\operatorname{L}\sqrt{\mathsf{D}_{\operatorname{last},2}}}{\sqrt{\mu}}\right)\tau_{2}^{2}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{3/2}}+\frac{\operatorname{L}^{1/2}\mathsf{D}_{\operatorname{last},4}^{2}\tau_{4}^{4}}{4\mu^{2}\mathsf{C}_{\operatorname{step},6}^{3/2}}\right)\end{split}$ (52)
@@
for any matrix $A\in\mathbb{R}^{d\times d}$, and $B_{2}\in\mathbb{R}^{d\times d}$ is a matrix, such that $\|B_{2}\|\leq\mathsf{C}_{b,2}$.
@@
$G(\theta)=\frac{1}{2}\left(\int_{0}^{1}t^{2}\nabla^{4}f(t\theta^{\star}+(1-t)\theta)\,dt\right)(\theta-\theta^{\star})^{\otimes 3}\,.$
@@
where $B\in\mathbb{R}^{d\times d}$ satisfies $\|B\|\leq\mathsf{C}_{b,3}$. Using (58) together with C 1($6$), we obtain that $\bar{\Sigma}_{\gamma}$ is a solution to the matrix equation
@@
$\|B^{\prime}\|_{{F}}\leq\mathsf{C}_{b,3}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{3/2}}+\frac{\operatorname{L}^{1/2}\mathsf{D}_{\operatorname{last},4}^{2}\tau_{4}^{4}}{4\mu^{2}\mathsf{C}_{\operatorname{step},6}^{3/2}}\,.$ (60)
@@
$\leq\frac{\sqrt{\gamma}\operatorname{L}\tau_{2}^{2}}{\mu}+\frac{1}{2}\left(\mathsf{C}_{b,3}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{3/2}}+\frac{\operatorname{L}^{1/2}\mathsf{D}_{\operatorname{last},4}^{2}\tau_{4}^{4}}{4\mu^{2}\mathsf{C}_{\operatorname{step},6}^{3/2}}\right)$
$\leq\frac{\sqrt{\operatorname{L}}\tau_{2}^{2}}{\sqrt{\mathsf{C}_{\operatorname{step},6}}\mu}+\frac{1}{2}\left(\mathsf{C}_{b,3}+\frac{\operatorname{L}\mathsf{D}_{\operatorname{last},6}^{3/2}\tau_{6}^{3}}{2\mu^{3/2}}+\frac{L^{1/2}\mathsf{D}_{\operatorname{last},4}^{2}\tau_{4}^{4}}{4\mu^{2}\mathsf{C}_{\operatorname{step},6}^{3/2}}\right)\,$,
@@
$\|\int_{\mathbb{R}^{d}}\mathcal{C}(\theta)\pi_{\gamma}(\mathrm{d}\theta)-\mathcal{C}(\theta^{\star})\|_{{F}}\leq\mathsf{C}_{b,3}\gamma^{1/2}\,$,
where the constant $\mathsf{C}_{b,3}$ is given by
$\mathsf{C}_{b,3}=\left(\frac{\operatorname{L}^{2}\mathsf{D}_{\operatorname{last},2}}{\mu^{3/2}}+\frac{\operatorname{L}\sqrt{\mathsf{D}_{\operatorname{last},2}}}{\sqrt{\mu}}\right)\tau_{2}^{2}\,.$ (63)